2 kg of ice at -20°C is mixed with 5 kg of water at 20°C in an insulating vessel having negligible heat capacity .cal final mass of water remaining in the container

Grade:12

1 Answers

Khimraj
5 years ago
The final mass of water in the container will be 6 kg
Explanation:
Heat lost by water when its temperature drops from 20°C to 0°C
= msΔt
= 5 x 1 x 20
= 100 kcal
Heat gain by water when its temperature rises from -20°C to 0°C
= msΔt
= 2 x 0.5 x 20 
= 20 kcal
Hence the remaining heat will be used to melt the ice into water
remaining heat = 100 - 20 = 80 kcal
given the latent heat of fusion of ice is 80 kcal/kg
let m kg of ice is converted to water, hence
m x latent heat = heat remained
=> m x 80 = 80
=> m = 1 kg
Hence the final mass of water will be 5+1 = 6 kg